# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v1

@@ 1 Introduction
    Preference inversion. Judge scores $S$ are treated as if they were rewards $Y$ on the same scale. They are not. Without calibration, higher $S$ can predict lower $Y$, meaning the optimization target is fundamentally misaligned with the evaluation target. The “You’re absolutely right!” phenomenon (sycophantic responses scoring high on early metrics but low on actual user value) is preference inversion in production.
@@
    Invalid confidence intervals. Naive CIs on uncalibrated judge scores achieve near-0% coverage: their 95% CIs capture the true value in 0/50 seeds. Calibration with Oracle-Uncertainty-Aware (OUA) inference dramatically improves coverage. (Figure 5: at 5% oracle fraction, calibration uncertainty contributes over half of total variance.)
@@
Recent work recognizes that LLM judges require calibration (Lee2025LLMJudgeReporting), but existing methods address only the simplest setting: single-model accuracy with binary outcomes. Guerdan2025DoublyRobustLLM use doubly robust methods to address external validity when LLM judges are prompted with demographic personas, a problem complementary to our focus on judge calibration for policy ranking. The harder problem of ranking multiple policies with continuous rewards under distribution shift remains inadequately addressed: no existing method provides calibration, overlap diagnostics, and valid uncertainty quantification in a unified framework. Without such a solution, practitioners either pay for full oracle labeling or accept invalid comparisons. CJE fills this gap.
@@
  • OUA inference: Delete-one-fold jackknife that propagates calibration uncertainty into confidence intervals, dramatically improving coverage.
These components instantiate a single design rule, Design-by-Projection (DbP), that encodes justified knowledge as model restrictions. Building on established semiparametric theory, we show that when justified knowledge defines a restricted statistical model, the efficiency bound in the restricted model is at most the bound in the baseline model (Theorem 2); with cross-fitting, our estimators attain the surrogate information bound.
Beyond variance control, CJE includes a mean transport test: for each target policy $\pi^{\prime}$ or time period, we test whether the mean residual $\mathbb{E}[Y-f(S,X)]$ is zero. This is a necessary condition for reusing a single calibration function across policies or over time; if it fails, surrogate-only evaluation is systematically biased. In our Arena benchmark, the base-trained calibration transports cleanly to clone, premium, and prompt-engineered policies, but fails for an adversarial “unhelpful” policy, where mean residuals reveal a 0.31-0.31 level shift.
@@
We validate CJE on a large-scale benchmark: 4,961 Chatbot Arena prompts, five LLM policies (including an adversarial “unhelpful” policy), 13 estimators, and GPT-5 as oracle. Key findings:
  • Calibration works. The Direct Method with two-stage calibration achieves 94% pairwise ranking accuracy on average (vs. 38% for SNIPS). Naive CIs on uncalibrated scores achieve 0% coverage; calibration with OUA improves coverage to 85–87% (Direct) and 95–96% (stacked-DR).
@@
  • DR does not dominate. We expected DR to outperform Direct by combining logged data with fresh draws. Instead, Direct slightly outperforms DR (94.3% vs. 94.1%): under low overlap, DR’s IPS component adds noise rather than information.
@@
    Theory: The CLE diagnostic provides a novel explanation for why high ESS is insufficient for IPS-style estimation, formalizing a precision floor under limited overlap. The Design-by-Projection framework applies established semiparametric principles to justify CJE’s methods.
@@ 2 Background and Setup
computed via teacher forcing (TF). The target is the counterfactual value $V(\pi^{\prime})=\mathbb{E}[Y(\pi^{\prime})]$. We use the sample-mean-one normalization (SNIPS) when helpful.
@@
IPS/SNIPS estimate $V(\pi^{\prime})$ by reweighting logged outcomes (HorvitzThompson1952; Hajek1964Rejective; LiChuLangfordSchapire2011; SwaminathanJoachims2015CRM). The direct method (DM) plugs in $g(x)=\sum_{a}\pi^{\prime}(a\mid x)\,\hat{m}(x,a)$. Doubly robust (DR) estimators combine IPS and DM (BangRobins2005) and, with sample–splitting and cross–fitting, admit $\sqrt{n}$ inference under the standard one–of–two $n^{-1/4}$ product–rate condition (Bickel1993; VaartWellner2000; Kosorok2008; JiangLi2016; Chernozhukov2018; vanDerLaanRose2011TL; KallusUehara2020DRL). Teacher forcing (TF) provides sequence–level propensities/ratios, so these forms apply to sequence policies in principle; however, TF reliability depends on API implementation details (see Limitations).
@@
IPS variance scales with $\mathbb{E}[W_{\pi^{\prime}}^{2}]$ and deteriorates under limited overlap (Crump2009LimitedOverlap). We monitor stability with the effective sample size (ESS),
@@
We also track tail behavior via diagnostics (Hill1975TailIndex; Liu2001MC; Owen2013MCTME). Common stabilizers include truncation/clipping (Ionides2008TruncatedIS), overlap weighting (LiMorganZaslavsky2018), balancing objectives (Kallus2018Balanced), and covariate–shift reweighting (Shimodaira2000CovShift; SugiyamaKrauledatMueller2007IWCV).
@@
Calibration enforces identities under $\pi_{0}$: outcome calibration de-biases $g(X)$; ratio calibration enforces $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}]=1$ and $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}h]=\mathbb{E}_{\pi^{\prime}}[h]$ for a test class $h$; orthogonal moments enable honest inference (Chernozhukov2018). Recent work gives projection-based IPS/DR with stability guarantees (vanDerLaanLinCaroneLuedtke2025StabIPW; vanDerLaanLuedtkeCarone2024DRCalibration). For DR, IF orthogonality renders small calibration error second order (Bickel1993; Chernozhukov2018; vanDerLaanRose2011TL).
@@
Isotonic regression is the Euclidean projection onto the cone of monotone functions (PAVA) (Ayer1955PAVA; Barlow1972); it avoids extrapolation and weakly reduces dispersion by majorization (HardyLittlewoodPolya1952; MarshallOlkinArnold2011). CJE exploits this via mean-preserving isotonic projections for both reward calibration and weight stabilization (Section 3).
@@
Automatic judges (LLM-as-judge or preference models) provide scalable scoring (Ouyang2022Instruct; BaiEtAl2022; Zheng2023LLMasJudge; Kim2024PrometheusJudge; KocmiFedermann2023) but are correlational and may drift (DietzEtAl2025LLMJudges). Viewing $S$ as a surrogate connects to surrogate endpoints and mediation (Prentice1989Surrogate; RobinsGreenland1992; FrangakisRubin2002PS; Pearl2012Mediation; VanderWeele2015Book). The surrogate paradox, wherein a policy improves $S$ and $S$ correlates with $Y$ yet the policy harms $Y$, can arise even with high correlation (VanderWeele2013Surrogates); our transport test (3) provides a necessary check. Under mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=\mu(S)$), calibrating $R=f(S)$ preserves $V(\pi^{\prime})=\mathbb{E}[f(S^{\pi^{\prime}})]$ and supplies a one-dimensional index that stabilizes weights.
@@
Treating learned $R=\hat{f}(S)$ as fixed understates uncertainty when the oracle slice is small. CJE addresses this via Oracle-Uncertainty-Aware (OUA) inference, which propagates calibration uncertainty into confidence intervals using a delete-one-fold jackknife (EfronStein1981; Bickel1993). For ensemble estimation, CJE stacks multiple estimators by minimizing IF covariance over the simplex (Wolpert1992StackedGeneralization; vanderLaan2007SuperLearner). Details in Section 3.
@@ 3 Methods
Figure 1: CJE pipeline overview. A small oracle slice (5–25%) provides expensive oracle labels to train a calibration model ($S\to Y$). The learned mapping is then applied to bulk evaluation data where oracle labels are unavailable, enabling policy evaluation at a fraction of the cost. (In experiments, the oracle is GPT-5; in production this would typically be human raters or downstream KPIs.)
@@ 3.2 Weight calibration (SIMCal-W: unit-mean ratios via OOF stacking of $S$-monotone candidates)
Let $W_{\pi^{\prime}}^{\mathrm{m1}}$ be the sample-mean-one baseline (SNIPS). For each fold $k$:
@@
Remark (transport view). In the continuous case the ideal component is $m^{\star}(s)=p_{S\mid\pi^{\prime}}(s)/p_{S\mid\pi_{0}}(s)$; SIMCal-W stacks increasing and decreasing $S$-monotone candidates to minimize IF variance, yielding approximately monotone weights when one direction dominates.
@@ 3.5 Oracle-Uncertainty-Aware (OUA) inference
Standard CIs treat the calibration function $\hat{f}$ as fixed, ignoring first-stage estimation error. When the oracle slice is small, this understates total variance and produces dramatically narrow intervals. In our experiments, naive CIs on uncalibrated scores achieve 0% coverage; calibration with OUA improves coverage to 85–96%.
@@
OUA adds $K$ refits of AutoCal-R (isotonic regression) and $K$ re-evaluations of the estimator. Since both are $O(n\log n)$, the overhead is modest relative to the cost of teacher forcing and response generation.
@@ 4 Theory: EIF, Design-by-Projection, and Efficiency
In plain terms: Every justified assumption we encode as a model restriction is “free variance reduction”: we shrink the tangent space without changing the estimand.
Remark (prior art). Theorem 2 formalizes a well-known principle in semiparametric efficiency theory (Newey1990Semiparametric; Bickel1993); our contribution is naming it and systematically applying it to surrogate-based LLM evaluation via the Design-by-Projection framework.
@@
Let $R=\hat{f}(Z(S))$ be AutoCal-R (monotone in $S$ or two-stage index; cross-fitted), and let $W^{\mathrm{m1}}_{\pi^{\prime}}\!\triangleq W_{\pi^{\prime}}/\bar{W}_{n}$ denote the sample-mean-one (SNIPS) ratios, where $\bar{W}_{n}=n^{-1}\sum_{i}W_{\pi^{\prime},i}$. Let $\hat{W}_{\pi^{\prime}}$ be SIMCal-W weights (OOF stack ++ variance guard $\rho\geq 1$). Then $\widehat{V}_{\mathrm{IPS}}=\tfrac{1}{n}\sum_{i}\hat{W}_{\pi^{\prime},i}R_{i}\to_{p}V(\pi^{\prime})$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection (see App. A.3), $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ and $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$; the reference implementation uses a simpler normalization that empirically yields similar ESS gains.
@@
i.e., DR-CPO attains the surrogate efficiency bound. (Efficiency attainment under the product-rate condition is a standard property of DR estimators (BangRobins2005; Chernozhukov2018); our contribution is verifying that the CJE construction satisfies these conditions.)
@@
Let $\{\widehat{V}^{(e)}\}_{e\in\mathcal{E}}$ be regular, asymptotically linear estimators with centered IFs $\{\phi^{(e)}\}$, and let $\hat{\alpha}\in\arg\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}$ the empirical IF covariance. If $\hat{\Sigma}\to\Sigma$ uniformly on $\Delta$, then the stacked estimator $\widehat{V}^{(\hat{\alpha})}$ is asymptotically linear with IF $\phi^{(\alpha^{\star})}$ and variance $\min_{\alpha\in\Delta}\alpha^{\top}\Sigma\,\alpha\leq\min_{e}\Sigma_{ee}$. An outer split leaves the limit unchanged. (This extends ensemble learning from prediction, where Super Learner (vanderLaan2007SuperLearner) combines algorithms to minimize cross-validated risk, to asymptotic efficiency: we combine estimators by minimizing the variance of their combined influence function.)

### v2

@@ 1 Introduction
    Preference inversion. Judge scores $S$ are treated as if they were rewards $Y$ on the same scale. They are not. Without calibration, higher $S$ can predict lower $Y$, inverting rankings entirely. The “You’re absolutely right!” phenomenon illustrates the cost: sycophantic responses score high on cheap proxies but low on actual user value. Teams know they should evaluate against long-term outcomes, but oracle labels are expensive, so they settle for proxies that may point the wrong direction.
@@
    Invalid confidence intervals. Naive CIs on uncalibrated judge scores achieve near-0% coverage: their 95% CIs capture the true value in 0/50 seeds. Bootstrap inference with bias-corrected estimation dramatically improves coverage to $\sim$}95% for Direct mode. (Figure 6: at 5% oracle fraction, calibration uncertainty contributes over half of total variance.)
@@
Recent work recognizes that LLM judges require calibration (Lee et al., 2025), but existing methods address only the simplest setting: single-model accuracy with binary outcomes. Guerdan et al. (2025) use doubly robust methods to address external validity when LLM judges are prompted with demographic personas, a problem complementary to our focus on judge calibration for policy ranking. The harder problem of ranking multiple policies with continuous rewards under distribution shift remains inadequately addressed: no existing method provides calibration, overlap diagnostics, and valid uncertainty quantification in a unified framework. Without such a solution, practitioners either pay for full oracle labeling or accept invalid comparisons. CJE fills this gap.
@@
  • OUA inference: Bootstrap with augmented inverse-propensity-weighted (AIPW) bias correction ($\hat{\theta}_{\mathrm{aug}}$) that propagates calibration uncertainty into confidence intervals, achieving near-nominal coverage.
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019), learning $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, and applies it to LLM evaluation with auditable transport diagnostics. These components instantiate a single design rule, Design-by-Projection (DbP), that encodes justified knowledge as model restrictions. Building on established semiparametric theory, we show that when justified knowledge defines a restricted statistical model, the efficiency bound in the restricted model is at most the bound in the baseline model (Theorem 2); with cross-fitting, our estimators attain the surrogate information bound.
Beyond variance control, CJE includes a mean transport test: for each target policy $\pi^{\prime}$ or time period, we test whether the mean residual $\mathbb{E}[Y-f(S,X)]$ is zero. This is a necessary condition for reusing a single calibration function across policies or over time; if it fails, surrogate-only evaluation is systematically biased. In our Arena benchmark, the base-trained calibration transports cleanly to clone, premium, and prompt-engineered policies, but fails for an adversarial “unhelpful” policy, where mean residuals reveal a 0.31-0.31 level shift. This test operationalizes a key epistemological shift: rather than assuming transportability holds globally (an untestable identification condition), we treat it as a falsifiable hypothesis testable for each policy. When the test fails, the failure is visible and actionable: either recalibrate with policy-specific data or refuse absolute value claims for that cell. This transforms a brittle identification assumption into a standardized audit.
@@
We validate CJE on a large-scale benchmark: 4,961 Chatbot Arena prompts, five LLM policies (including an adversarial “unhelpful” policy), 13 estimators, and gpt-5-2025-08-07 as oracle. Key findings:
  • Calibration works. The Direct Method with two-stage calibration achieves 94% pairwise ranking accuracy on average (vs. 38% for SNIPS). Naive CIs on uncalibrated scores achieve 0% coverage; bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$ improves coverage to $\sim$}95% (Direct and stacked doubly-robust).
@@
  • Doubly-robust (DR) does not dominate. We expected DR to outperform Direct by combining logged data with fresh draws. Instead, Direct slightly outperforms DR (94.4% vs. 94.1%): under low overlap, DR’s IPS component adds noise rather than information.
@@
    Theory: CJE adapts the surrogate index framework to LLM evaluation with a key innovation: transportability becomes auditable rather than assumed via the mean transport test. Inference is grounded in the efficient influence function for missing-outcome estimation, explaining why bootstrap with $\hat{\theta}_{\mathrm{aug}}$ achieves $\sim$}95% coverage where naive methods achieve 0%. The CLE diagnostic provides a novel explanation for OPE failure despite high ESS, formalizing a precision floor under limited overlap.
@@ 2 Background and Setup
computed via teacher forcing (TF). The target is the counterfactual value $V(\pi^{\prime})=\mathbb{E}[Y(\pi^{\prime})]$. We use self-normalized importance sampling (SNIPS), which produces sample-mean-one weights, when helpful.
@@
CJE as surrogate index for LLM evaluation.
Conceptually, CJE adopts the surrogate index move (Athey et al., 2019): learn a mapping $f(S,X)\approx\mathbb{E}[Y\mid S,X]$ from cheap proxies to expensive outcomes, then use it for downstream estimation. For policy value estimation, we require only mean transport: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$, which we treat as auditable via a small oracle slice in each target context (3). We additionally impose monotonicity and mean-preservation as Design-by-Projection restrictions. These are not identification assumptions, but deliberate regularization choices that stabilize calibration and prevent preference inversions.
Athey et al.’s framework addresses cross-sample comparability when surrogates are learned in one environment and applied in another. CJE faces two analogous cross-environment challenges: (i) calibration transport (does $f$ learned under $\pi_{0}$ remain mean-unbiased under $\pi^{\prime}$?), addressed by the mean residual test; and (ii) action-space coverage (does the logger visit regions where $\pi^{\prime}$ concentrates?), addressed by CLE/TTC diagnostics. The key difference: we budget oracle labels in each target context to audit transport rather than assume it, making CJE an evaluation protocol with explicit failure detection rather than a brittle identification claim.
@@
IPS/SNIPS estimate $V(\pi^{\prime})$ by reweighting logged outcomes (Horvitz & Thompson, 1952; Hájek, 1964; Li et al., 2011; Swaminathan & Joachims, 2015). The direct method (DM) plugs in $g(x)=\sum_{a}\pi^{\prime}(a\mid x)\,\hat{m}(x,a)$. Doubly robust (DR) estimators combine IPS and DM (Bang & Robins, 2005) and, with sample–splitting and cross–fitting, admit $\sqrt{n}$ inference under the standard one–of–two $n^{-1/4}$ product–rate condition (Bickel et al., 1993; van der Vaart & Wellner, 2000; Kosorok, 2008; Jiang & Li, 2016; Chernozhukov et al., 2018; van der Laan & Rose, 2011; Kallus & Uehara, 2020). Teacher forcing (TF) provides sequence–level propensities/ratios, so these forms apply to sequence policies in principle; however, TF reliability depends on API implementation details (see Limitations).
@@
IPS variance scales with $\mathbb{E}[W_{\pi^{\prime}}^{2}]$ and deteriorates under limited overlap (Crump et al., 2009). We monitor stability with the effective sample size (ESS),
@@
We also track tail behavior via diagnostics (Hill, 1975; Liu, 2001; Owen, 2013). Common stabilizers include truncation/clipping (Ionides, 2008), overlap weighting (Li et al., 2018), balancing objectives (Kallus, 2018), and covariate–shift reweighting (Shimodaira, 2000; Sugiyama et al., 2007).
@@
Calibration enforces identities under $\pi_{0}$: outcome calibration de-biases $g(X)$; ratio calibration enforces $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}]=1$ and $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}h]=\mathbb{E}_{\pi^{\prime}}[h]$ for a test class $h$; orthogonal moments enable honest inference (Chernozhukov et al., 2018). Recent work gives projection-based IPS/DR with stability guarantees (van der Laan et al., 2025; 2024). For DR, IF orthogonality renders small calibration error second order (Bickel et al., 1993; Chernozhukov et al., 2018; van der Laan & Rose, 2011).
@@
Isotonic regression is the Euclidean projection onto the cone of monotone functions (PAVA) (Ayer et al., 1955; Barlow et al., 1972); it avoids extrapolation and weakly reduces dispersion by majorization (Hardy et al., 1952; Marshall et al., 2011). CJE exploits this via mean-preserving isotonic projections for both reward calibration and weight stabilization (Section 3).
@@
Automatic judges (LLM-as-judge or preference models) provide scalable scoring (Ouyang et al., 2022; Bai et al., 2022; Zheng et al., 2023; Kim et al., 2024; Kocmi & Federmann, 2023) but are correlational and may drift (Dietz et al., 2025). Viewing $S$ as a surrogate connects to surrogate endpoints and mediation (Prentice, 1989; Robins & Greenland, 1992; Frangakis & Rubin, 2002; Pearl, 2012; VanderWeele, 2015). The surrogate paradox, wherein a policy improves $S$ and $S$ correlates with $Y$ yet the policy harms $Y$, can arise even with high correlation (VanderWeele, 2013); our transport test (3) provides a necessary check. Under mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=\mu(S)$), calibrating $R=f(S)$ preserves $V(\pi^{\prime})=\mathbb{E}[f(S^{\pi^{\prime}})]$ and supplies a one-dimensional index that stabilizes weights.
[…]
Figure 1: Judge-as-surrogate under policy and environment shift. $X$: prompt; AA: response; $S$: judge score; $Y$: oracle label; $\pi$: policy; $E$: environment. $S$ and $Y$ are parallel measurements of the response. Solid arrows: causal structure ($\pi\to A$: policy determines response). Dashed red arrows: mechanism-shift edges indicating the $S$-$Y$ calibration may differ across policies or environments; these are not causal effects of $\pi$ or $E$ on $Y$ holding $(X,A)$ fixed. CJE tests whether calibration transports via $\mathbb{E}[Y-f(S,X)]=0$.
Surrogate interpretation.
We use the judge score $S$ as a surrogate measurement for the expensive outcome $Y$, not as a causal mediator. Given a calibration function $f$ learned on an oracle slice, define the residual $\varepsilon:=Y-f(S,X)$. For any target policy $\pi^{\prime}$, the value estimation error decomposes exactly as
$\mathbb{E}_{\pi^{\prime}}[Y]-\mathbb{E}_{\pi^{\prime}}[f(S,X)]\;=\;\mathbb{E}_{\pi^{\prime}}[\varepsilon].$ (2)
Thus, surrogate-only evaluation is mean-unbiased for policy $\pi^{\prime}$ if and only if the mean residual transports: $\mathbb{E}_{\pi^{\prime}}[\varepsilon]=0$. We treat this as an auditable assumption rather than an identification axiom: we estimate and test $\mathbb{E}_{\pi^{\prime}}[\varepsilon]$ on a small oracle audit slice per policy (or time window), and re-anchor calibration (or refuse level estimates) when the test fails (3).
This mean-transport condition is strictly weaker than Prentice-style surrogacy, which requires $Y\perp\pi^{\prime}\mid(S,X)$; it suffices for unbiased mean value estimation but does not guarantee conditional calibration within subgroups or at distribution tails. When finer guarantees are needed, the audit can be strengthened from a single mean to decile-binned residual checks, detecting compensating errors that cancel in the overall mean.
@@
Treating learned $R=\hat{f}(S)$ as fixed understates uncertainty when the oracle slice is small. CJE addresses this via Oracle-Uncertainty-Aware (OUA) inference, which propagates calibration uncertainty into confidence intervals using a delete-one-fold jackknife (Efron & Stein, 1981; Bickel et al., 1993). For ensemble estimation, CJE stacks multiple estimators by minimizing IF covariance over the simplex (Wolpert, 1992; van der Laan et al., 2007). Details in Section 3.
@@ 3 Methods
Figure 2: CJE pipeline overview. A small oracle slice (5–25%) provides expensive oracle labels to train a calibration model ($S\to Y$). The learned mapping is then applied to bulk evaluation data where oracle labels are unavailable, enabling policy evaluation at a fraction of the cost. (In experiments, the oracle is gpt-5-2025-08-07; in production this would typically be human raters or downstream KPIs.)
@@ 3.2 Weight calibration (SIMCal-W: unit-mean ratios via OOF stacking of $S$-monotone candidates)
Let $W_{\pi^{\prime}}^{\mathrm{m1}}$ be the self-normalized (SNIPS) baseline. For each fold $k$:
@@
Remark (transport view). In the continuous case the ideal component is $m^{\star}(s)=p_{S\mid\pi^{\prime}}(s)/p_{S\mid\pi_{0}}(s)$; SIMCal-W stacks increasing and decreasing $S$-monotone candidates to minimize IF variance, yielding approximately monotone weights when one direction dominates. When $m^{\star}(S)$ is not approximately monotone, isotonic projection trades bias for variance; in the low-overlap regimes where this matters most, we recommend Direct over IPS/DR (see CLE diagnostic, Section 2).
@@ 3.5 Oracle-Uncertainty-Aware (OUA) inference
Standard CIs treat the calibration function $\hat{f}$ as fixed, ignoring first-stage estimation error. When the oracle slice is small, this understates total variance and produces dramatically narrow intervals. In our experiments, naive CIs on uncalibrated scores achieve 0% coverage; bootstrap inference with bias-corrected estimation achieves $\sim$}95% coverage for Direct mode.
@@
Bootstrap with bias-corrected estimation.
While the jackknife captures variance, it does not correct for bias in the plug-in estimator $\widehat{V}=n^{-1}\sum_{i}\hat{f}(S_{i})$. This bias arises because $\hat{f}$ is fit on the oracle slice and applied to all samples, creating a covariance between the calibration function and the evaluation samples. We address this with the one-step estimator derived from the efficient influence function for missing-outcome estimation (Section C.13):
$\hat{\theta}_{\mathrm{aug}}=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(S_{i})+\frac{1}{|L|}\sum_{i\in L}\big(Y_{i}-\hat{f}^{(-i)}(S_{i})\big),$ (3)
where $\hat{f}$ is the full calibrator, $\hat{f}^{(-i)}$ is the calibrator trained on folds excluding sample $i$’s cluster, and $L$ is the oracle slice. The first term uses the full calibrator for lower variance; the second term uses cluster-out-of-fold predictions to avoid overfitting in the residual correction. If oracle sampling is non-uniform (e.g., oversampling uncertain prompts), replace the constant $|L|/n$ with sample-specific inclusion probabilities $e(Z_{i})$ in the general AIPW form (Section C.13). Combined with bootstrap (refitting the calibrator in each replicate), this achieves $\sim$}95% coverage across oracle fractions (Table 3), compared to 70–89% for jackknife-only OUA.
@@
Bootstrap with $\hat{\theta}_{\mathrm{aug}}$ requires $B$ refits of AutoCal-R (we use B=2, [^0] B{=}2{,}000). Since isotonic regression is $O(n\log n)$, the overhead remains modest relative to the cost of teacher forcing and response generation. The full bootstrap procedure (Algorithm 7) uses prompt-level cluster resampling to preserve within-prompt dependence (Cameron et al., 2008), with a resample-until-valid loop ensuring each replicate has sufficient oracle labels.
@@ 4 Theory: EIF, Design-by-Projection, and Efficiency
In plain terms: Every justified assumption we encode as a model restriction is “free variance reduction”: we shrink the tangent space without changing the estimand. If a restriction is violated, projection introduces bias; our diagnostics and gates (Sections 2 and 6) detect the regimes where this matters.
Remark (prior art). Theorem 2 formalizes a well-known principle in semiparametric efficiency theory (Newey, 1990; Bickel et al., 1993); our contribution is naming it and systematically applying it to surrogate-based LLM evaluation via the Design-by-Projection framework.
@@
Let $R=\hat{f}(Z(S))$ be AutoCal-R (monotone in $S$ or two-stage index; cross-fitted), and let $W^{\mathrm{m1}}_{\pi^{\prime}}\!\triangleq W_{\pi^{\prime}}/\bar{W}_{n}$ denote the self-normalized (SNIPS) ratios, where $\bar{W}_{n}=n^{-1}\sum_{i}W_{\pi^{\prime},i}$. Let $\hat{W}_{\pi^{\prime}}$ be SIMCal-W weights (OOF stack ++ variance guard $\rho\geq 1$). Then $\widehat{V}_{\mathrm{IPS}}=\tfrac{1}{n}\sum_{i}\hat{W}_{\pi^{\prime},i}R_{i}\to_{p}V(\pi^{\prime})$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection (see App. A.3), $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ and $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$; the reference implementation uses a simpler normalization that empirically yields similar ESS gains.
@@
i.e., DR-CPO attains the surrogate efficiency bound. (Efficiency attainment under the product-rate condition is a standard property of DR estimators (Bang & Robins, 2005; Chernozhukov et al., 2018); our contribution is verifying that the CJE construction satisfies these conditions.)
@@
Let $\{\widehat{V}^{(e)}\}_{e\in\mathcal{E}}$ be regular, asymptotically linear estimators with centered IFs $\{\phi^{(e)}\}$, and let $\hat{\alpha}\in\arg\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}$ the empirical IF covariance. If $\hat{\Sigma}\to\Sigma$ uniformly on $\Delta$, then the stacked estimator $\widehat{V}^{(\hat{\alpha})}$ is asymptotically linear with IF $\phi^{(\alpha^{\star})}$ and variance $\min_{\alpha\in\Delta}\alpha^{\top}\Sigma\,\alpha\leq\min_{e}\Sigma_{ee}$. An outer split leaves the limit unchanged. (This extends ensemble learning from prediction, where Super Learner (van der Laan et al., 2007) combines algorithms to minimize cross-validated risk, to asymptotic efficiency: we combine estimators by minimizing the variance of their combined influence function.)